OpenStudy (anonymous):

Solve the linear system by using substitution or elimination. 3x + 2y = 4 -6x - 4y = -8

OpenStudy (solomonzelman):

Multiply the 1st equation by 2 and add, tell me what you get...

OpenStudy (solomonzelman):

3x + 2y = 4 ---> times 2 ----> 6x+4y=8 -6x - 4y = -8 6x+4y=8 and -6x - 4y = -8 you can see that the second equation is the ' first times -1 ' there could be therefore, infinite number of solutions.

OpenStudy (solomonzelman):

when you add them you just get 0=0